# Towards Intent-based User Interfaces: Charting the Design Space of Intent-AI Interactions Across Task Types

### v1

@@ 1. Introduction
Human-computer interaction (HCI) involves humans expressing intent for a task, which the machine then interprets, executes, and communicates the results back to the user. Historically, the capabilities of machines to interpret human intents and execute tasks were limited, resulting in complex and burdensome processes for expressing these intents. This complexity often necessitated the mastery of intricate programming languages for command-line interfaces or the manipulation of buttons, dropdown menus, and other elements within graphical user interfaces (GUIs). The advent of generative AI, particularly the emergence of large language models, enables humans — at least for simple tasks like text summarization — to convey their intents to machines using natural language prompts, which are then translated into actions. Building on these shifts in machine capabilities and interaction paradigms, we are inspired to explore the concept of Intent-based User Interfaces (IUIs). Here, ”intent” refers to the overarching purpose or objective at a higher level in the causal chain of a task, rather than the specific steps such as direct manipulations on complex GUIs. For instance, in a data analysis programming environment, the primary intent is to understand trends in the data, as opposed to manually inputting column names in plotting functions. This leads us to our research question for the design space of IUI: What kinds of intent expression, iteration and organization are needed for different types of tasks?
@@ 2.3. Complex and Interdependent Tasks
Beyond the atomic creative tasks are complex and interdependent tasks such as active search (Palani et al., 2021b, a), data visualization (Vaithilingam et al., 2024), video editing (Wang et al., 2024), active search (Palani et al., 2021a, b), neurocognitive disorders screening (Ding et al., 2022), and storytelling in text (Yuan et al., 2022; Singh et al., 2022) and images (Yan et al., 2023). These tasks demand not only domain expertise but also capabilities for planning, reasoning, ideating, and maintaining context over extended periods to create coherent and innovative content. Researchers have developed LLM-supported tools like LAVE (Wang et al., 2024) for video editing, DynaVis (Vaithilingam et al., 2024) for data visualization, and XCreation (Yan et al., 2023) for cross-modal storytelling to facilitate these intricate interactions.
@@ 3. Study Plan and Preliminary Results
Our research comprises three studies, each examining intent-AI interactions across tasks varying in creativity and complexity. The initial study delves into fixed-scope content curation tasks, utilizing news headline generation as the investigative lens. The first study examines fixed-scope content curation tasks, employing news headline generation as the medium for exploration. The second study transitions to probing atomic creative tasks, with a focus on analogy generation to uncover underlying creative processes. Lastly, the third ongoing study explores the complex and interdependent tasks associated with exploratory visual data analysis. We hope to use these three studies to collectively offer a comprehensive understanding of intent-AI interaction across a spectrum of task.
@@ 3.2. Study 2 on Cross-Domain Analogy Generation: Intent Iteration for Atomic Creative Tasks
To understand how users engage with this innovative interface, we conducted a user study involving 10 data analysts, ranging from novices to experienced professionals, complemented by semi-structured interviews. Our findings reveal differences in interaction patterns between the two groups. Experienced data analysts predominantly engaged in iterative refinement of their visualizations, often using detailed, command-like expressions of intent, such as ”Correlation plot of data - (+) correlation coefficient in red, (-) correlation coefficient in blue; include all variables that are continuous”. This approach yielded high-quality outputs but limited the exploration of alternative analytical perspectives, potentially overlooking the exploratory nature of the task.

### v2

@@ 1. Introduction
Human-computer interaction (HCI) involves humans expressing intent for a task, which the machine then interprets, executes, and communicates the results back to the user. Historically, the capabilities of machines to interpret human intents and execute tasks were limited, resulting in complex and burdensome processes for expressing these intents. This complexity often necessitated the mastery of intricate programming languages for command-line interfaces or the manipulation of buttons, dropdown menus, and other elements within graphical user interfaces (GUIs). The advent of generative AI, particularly the emergence of large language models, enables humans — at least for simple tasks like text summarization — to convey their intents to machines using natural language prompts, which are then translated into actions. Building on these shifts in machine capabilities and interaction paradigms, we are inspired to explore the concept of Intent-based User Interfaces (IUIs). Here, ”intent” refers to the overarching purpose or objective at a higher level in the causal chain of a task, rather than the specific steps such as direct manipulations on complex GUIs. For instance, in a data analysis programming environment, the primary intent is to understand trends in the data, as opposed to manually inputting column names in plotting functions. This leads us to our research question for the design space of IUI: What kinds of intent-AI interaction are needed across different types of tasks?
@@ 2.3. Complex and Interdependent Tasks
Beyond the atomic creative tasks are complex and interdependent tasks such as active search (Palani et al., 2021b, a), data visualization (Vaithilingam et al., 2024), video editing (Wang et al., 2024), active search (Palani et al., 2021a, b), design workshops (MacNeil et al., 2023), neurocognitive disorders screening (Ding et al., 2022), and storytelling in text (Yuan et al., 2022; Singh et al., 2022) and images (Yan et al., 2023). These tasks demand not only domain expertise but also capabilities for planning, reasoning, ideating, and maintaining context over extended periods to create coherent and innovative content. Researchers have developed LLM-supported tools like LAVE (Wang et al., 2024) for video editing, DynaVis (Vaithilingam et al., 2024) for data visualization, and XCreation (Yan et al., 2023) for cross-modal storytelling to facilitate these intricate interactions.
@@ 3. Study Plan and Preliminary Results
Our research comprises three studies, each examining intent-AI interactions across tasks varying in creativity and complexity. The first study examines fixed-scope content curation tasks, employing news headline generation as the medium for exploration. The second study transitions to probing atomic creative tasks, with a focus on analogy generation to uncover underlying creative processes. Lastly, the third ongoing study explores the complex and interdependent tasks associated with exploratory visual data analysis. We hope to use these three studies to collectively offer a comprehensive understanding of intent-AI interaction across a spectrum of task.
@@ 3.2. Study 2 on Cross-Domain Analogy Generation: Intent Iteration for Atomic Creative Tasks
To understand how users engage with this experimental interface, we conducted a user study involving 10 data analysts, ranging from novices to experienced professionals, complemented by semi-structured interviews. Initial findings revealed differences in interaction patterns between the two groups. Experienced data analysts predominantly engaged in iterative refinement of their visualizations, often using detailed, command-like expressions of intent, such as ”Correlation plot of data - (+) correlation coefficient in red, (-) correlation coefficient in blue; include all variables that are continuous”. This approach yielded high-quality outputs but limited the exploration of alternative analytical perspectives, potentially overlooking the exploratory nature of the task.